Thomas uses the wall tiles behind the canvas to estimate it's dimensions. If each square tile is 6 inches wide, what are the dimensions of the canvas?
1
Expert's answer
2020-10-13T19:15:15-0400

Let suppose n number of tiles is required to covered the whole canvas,


Area of 1 tile="(side)^2"

="6^2"

=36 square units


so Area of canvas ="n\times" Area of 1 tile

=36n square units

Case:- If the canvas is of square shape the side of canvas

="\sqrt{area of canvas}"

="\sqrt{36n}"

=6"\sqrt{n}" units